OpenStudy (anonymous):

mother bought 2 1/2 pounds of berries on monday. Sammy eat 1/4 of the berries. how many pounds of berries sammy eat on monday. I wrote: she ate 0.625 on monday but not sure and need help with 2nd and 3rd part. Thank you for helping Second part: how many pounds leftover? third part: on tuesday sammy eat 1/3 of the leftovers. hom many pounds did sammy eat on tuesday? Thank you

OpenStudy (anonymous):

didn't you just post this problem beforE?

OpenStudy (lilg132):

0.625 you did post this question before hence why i still remember the answer

OpenStudy (anonymous):

for the first part just multiply 2 1/2 times 1/4. what do you get? it's probably easiest to put 2 1/2 into an improper fraction first.

OpenStudy (lilg132):

or change it to numbers and then do it

OpenStudy (anonymous):

i got the 0.625.. but what do i do with the 2nd and 3rd part?

OpenStudy (anonymous):

0.625 is not correct for part a.

OpenStudy (anonymous):

5/2 times 1/4 = 5/8

OpenStudy (anonymous):

oh! ha, you're right, sorry, I was computing part b. So part b is asking how many berries are left over?

OpenStudy (anonymous):

if you start with 2 1/2, and eat 5/8, what's left?

OpenStudy (anonymous):

ok so.. 2 1/2 = 5/2 minus 5/8?

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

15/8

OpenStudy (anonymous):

I'd first convert 5/2 into something/8.

OpenStudy (anonymous):

yep, 20/8 minus 5/8 is 15/8. nice!

OpenStudy (anonymous):

now he has 15/8 and he eats 1/3 of that. so, multiply.

OpenStudy (anonymous):

15/24?

OpenStudy (anonymous):

you got it! you can reduce 15/24 now, since there's a common factor of 3.

OpenStudy (anonymous):

5/8

OpenStudy (anonymous):

that's it!

OpenStudy (anonymous):

thank you..!!

OpenStudy (anonymous):

no problem